### v2

# A Certifiable Machine Learning-Based Pipeline to Predict Fatigue Life of Aircraft Structures

Abstract
Fatigue life prediction is essential in both the design and operational phases of any aircraft, and in this sense safety in the aerospace industry requires early detection of fatigue cracks to prevent in-flight failures. Robust and precise fatigue life predictors are thus essential to ensure safety. Traditional engineering methods, while reliable, are time consuming and involve complex workflows, including steps such as conducting several Finite Element Method (FEM) simulations, deriving the expected loading spectrum, and applying cycle counting techniques like peak-valley or rainflow counting. These steps often require collaboration between multiple teams and tools, added to the computational time and effort required to achieve fatigue life predictions. Machine learning (ML) offers a promising complement to traditional fatigue life estimation methods, enabling faster iterations and generalization, providing quick estimates that guide decisions alongside conventional simulations. In this paper, we present a ML-based pipeline that aims to estimate the fatigue life of different aircraft wing locations given the flight parameters of the different missions that the aircraft will be operating throughout its operational life. We validate the pipeline in a realistic use case of fatigue life estimation, yielding accurate predictions alongside a thorough statistical validation and uncertainty quantification. Our pipeline constitutes a complement to traditional methodologies by reducing the amount of costly simulations and, thereby, lowering the required computational and human resources.
keywords:
Fatigue Life , Machine Learning , Aerospace Safety , Aerospace Certification , Statistical validation
journal: International Journal of Fatigue
[…]
[1]organization=ETSIAE-UPM - School of Aeronautics, Universidad Politécnica de Madrid,addressline=Plaza Cardenal Cisneros 3, postcode=E-28040, city=Madrid, country=Spain
[…]
[2]organization=Airbus Defence & Space, Fatigue & Damage Tolerance,city=Getafe, country=Spain
[…]
[3]organization=Institute for Cross-Disciplinary Physics and Complex Systems (IFISC), CSIC-UIB,city=Palma de Mallorca, country=Spain
[…]
[4]organization=Center for Computational Simulation, Universidad Politécnica de Madrid, Campus de Montegancedo, Boadilla del Monte,postcode=28660, city=Madrid, country=Spain
List of Acronyms
AD Anderson–Darling
CFD Computational Fluid Dynamics
CI95 95% Confidence Interval
CLT Central Limit Theorem
CV Coefficient of Variation
EASA European Aviation Safety Agency
ESDU Engineering Sciences Data Unit
FEM Finite Element Method
GAG Ground Air Ground
G&M Gusts and Maneuvers
IQR Interquartile Range
KS Kolmogorov–Smirnov
MAE Mean Absolute Error
ML Machine Learning
MLP Multilayer Perceptron
MRE Mean Relative Error
PSE Principal Structural Element
PSD Power Spectral Density
1 Introduction
Damage estimation caused by fatigue [grover1966fatigue] is a critical aspect of aeronautical engineering, as failure to predict it accurately can lead to catastrophic structural breakdowns and compromise both flight safety and operational efficiency. Likewise, fatigue damage assessment is essential for implementing predictive maintenance strategies that extend the life of aircraft and reduce unplanned downtime. Standard fatigue prediction methods rely on a convoluted and resource-intensive process that combines service history of aircraft of similar structural design with aerodynamic and Finite Element Methods simulations and cumulative damage modeling to guide expert-based decision making (CS-25.571 in [easaCS25]). The complexity of the whole process makes the development of tailored maintenance routines for new aircraft operators particularly expensive and time-consuming, involving multidisciplinary teams and a range of iterative procedures.
Interestingly, the advent of Machine Learning (ML) techniques and the rapid incorporation of these tools within the aeronautical engineering community (encompassing fields from aerodynamics [linse1993identification, suresh2003lift, secco2017artificial], stress prediction [lacasa2025certification] or aeroacoustics [ALGUACIL2021116285, ruettgers2020acoustic] to airflow optimization [ramos2025transfer] –see [LeClainche2023] for a recent review–) provides a realistic opportunity to streamline and optimize the procedures for fatigue-related damage prediction, as these techniques potentially offer near-instantaneous inference of fatigue life, once properly trained. Specifically in fatigue-related problems, ML has been recently used for load spectrum reconstruction [graziani2024prediction, Valdes2017], fatigue life prediction [palczynski2022application, chen2023pinn], and crack growth modeling [Zhang2020, Younis2023]. However, most approaches treat these steps separately and rely on hybrid frameworks that combine ML with traditional tools, lacking end-to-end ML-enabled fatigue life prediction pipelines.
At the same time, there is an urge of incorporating fully certifiable ML-based procedures within the aeronautical industry. As a matter of fact, a key consideration for ML in safety-critical applications is the ability to certify the predictive model. Aeronautical authorities such as the European Aviation Safety Agency (EASA) have outlined the main challenges and goals of these industries for trustworthy ML adoption [CoDANN, CoDANN2], and recent research has started to propose practical methodologies for statistical validation of ML pipelines in industrial settings [lacasa2025certification], encompassing data preparation, model training and interpretability, and multiscale uncertainty quantification.
This work proposes a fully integrated and certifiable ML-based pipeline capable of predicting fatigue life at various locations of an aircraft wing, based solely on flight parameters. The pipeline mostly uses a suite of different multilayer perceptrons (MLPs) to approximate different input-output relationships within the fatigue damage assessment. An MLP is an archetypal method in supervised machine learning [goodfellow2016deep] used to infer (regress or classify) complex input-output representations $x\to y$, where $x\in\mathbb{R}^{m}$ and $y$ is usually another vector for regression tasks or an element from a discrete set in classification tasks. Often visualized as an (artificial neural) network, an MLP consists in a (deep) feedforward neural network representing an overparameterized nonlinear function $\mathscr{F}(x;{\Omega})$, where $\Omega=\{w_{k}\}$ –the set of trainable parameters– identify with the weights of the links between neurons (see [goodfellow2016deep] for details). The precise architecture of the MLP (number of neurons per layer, number of layers) as well as the precise optimization scheme to train the network are defined in a case-by-case basis.
Our approach enables end-to-end learning of the physical process –from stress estimation to load cycle characterization and fatigue damage inference– reducing the need for intermediate, time consuming CFD or FEM simulations. At the same time, the pipeline enforces a statistically solid validation protocol which, overall, aims to pave the way for certifiable ML-driven fatigue assessment. Using real fatigue data from 38 wing locations of a tactical aircraft for different operational profiles, we show that the pipeline provides accurate fatigue life prediction along with a thorough characterization of the model’s uncertainty.
The rest of this paper is structured as follows: in Section 2, we present the full methodology: we define the physical problem under study, the dataset, the set of ML techniques and software tools, and the statistical techniques used for a thorough validation. Then, in Section 3 we present the results, which we anticipate yield a successful ML-driven fatigue life prediction tool. Finally, conclusions and future directions are discussed in Section 4.
2 Methodology
In this section, we present the methodology adopted in this work. First, in subsections 2.1 and 2.2, we describe the physical problem being modeled using machine learning techniques and explain how it is decomposed into smaller, learnable subproblems. Then, in subsection 2.3, we detail the available dataset and the data-splitting strategy used for training, which plays a crucial role in the performance of the models. Finally, in subsection 2.4, we add a few comments on the ML and statistical software libraries we have used.
[…]
2.1 Breaking down the fatigue damage assessment
The goal of this study is to develop a machine learning pipeline to predict fatigue life (expressed as the number of flights to failure) at multiple wing locations –known as principal structural elements (PSEs)–, when flying a specific mission (i.e., a flight characterized by specific flight parameters and operational objectives). In the traditional life estimation method we aim to model, all PSEs are considered non-damage-tolerant, and consequently failure is defined as the number of flights until a detectable crack initiates. Here we adopt this very same definition. In what follows we unfold the process and breakdown the whole supervised learning task in a sequence of interpretable subtasks, see Fig. 1 for an illustration of the different steps.
Figure 2: A mission flight plan, partitioned as a sequence of segments. Each segment is parametrized by a 12-dimensional feature vector $\bf P$ whose elements include variables such as the degree of extension of the flaps in that segment, the true airspeed during the segment, etc (see Table 8 for details.)
Mission segmentation – The assessment process starts by defining the mission plan (subparagraph 6b in CS-25.571, [easaCS25]). This process consists in partitioning a mission into a structured sequence of flight segments, see Fig. 2. Each of these segments characterize different specific phases of the flight, so that the categorical variable segment $\in$ {taxi, climb, cruise, descent, approach}. It is also useful to classify segments in the ground phase (taxi) or in the flight phase (every other segment except taxi) [^1] While in rigor there exist ground segments other than taxi (e.g. take-off, braking, rotation, etc), these are not explicitly taken into account in the models as the wing PSEs don’t show any significant damage induced by these type of segments.. Accordingly, a mission flight plan is simply represented as a sequence of segments. At the same time, each segment is characterized by a vector of 12 flight parameters ${\bf P}\in\mathbb{R}^{12}$, that includes variables such as the degree of extension of the flaps in that segment, the true airspeed during the segment, the altitude at which the segment takes place, and so on (see Table 8 for details). Accordingly, within each mission, any given flight is described in terms of a sequence of segment feature vectors $\{{\bf P}_{i}\}$ with class metadata (segment class and ground/air typification).
From aerodynamic loads to PSE stresses – Once the mission has been segmented and parameterized, Computational Fluid Dynamics (CFD) simulations are performed in each mission segment to obtain the corresponding aerodynamic loads. These loads are then propagated onto the structural level using Finite Element Method (FEM) simulations, so as to obtain the stresses at the different principal structural elements (PSEs) of interest (appendix 2 in CS-25.571, [easaCS25]). The resulting values will be used for the stress spectrum generation. For each PSE, we have a particular four-dimensional stress vector ${\bf S}\in\mathbb{R}^{4}$, whose elements are ${\bf S}=(1g,\Delta\text{vman},\Delta\text{vgust},\Delta{\text{turn}})$, where $1g$ is the mean equilibrium stress at the studied PSE in steady flight (i.e. the structural response under nominal flight conditions), whereas the other three variables are incremental stresses at the studied PSE caused by a standard load factor of $1.5g$ due to vertical maneuvers, vertical gusts and turns. These latter variables capture the effects of atmospheric turbulence and pilot-induced maneuvers (see Table 9 for details). The incremental stresses values depend on the modeling. To account for gusts, Engineering Sciences Data Unit (ESDU) data or Power Spectral Density (PSD) models (CS-25.341 in [easaCS25]) can be used, and for maneuvers, military regulations specify the requirements [mil8861b]. This step is computationally intensive and must be repeated for each mission segment individually, making it one of the most resource-consuming parts of the process.
From stresses to GAG and G&M cycles – With the stress data obtained in the previous step and historical flight records from a similar aircraft, two types of load cycles are derived for each mission: the so-called Ground-Air-Ground (GAG) cycle and the cycles induced by gusts and maneuvers (G&M). The GAG cycle represents the dominant stress fluctuation experienced as the aircraft transitions from ground to flight segments. This cycle typically involves the largest stress amplitude of the flight (minimum to maximum stress) and occurs once per flight. In contrast, gusts and maneuvers generate multiple smaller cycles within each flight segment. These are constructed using statistical load distributions and serve as an approximation of the fatigue loads that the aircraft will encounter across all the mission flights. Each resulting cycle is represented as a triplet: the maximum stress, the minimum stress, and the number of times that specific cycle occurs. Together, the cycles induced by gusts and maneuvers, along with the GAG cycle (derived from them) constitute the two main sources of fatigue damage—typically with the GAG cycle being the dominant contributor.
Final assessment of fatigue life – Once the load cycles have been defined, the fatigue damage and corresponding life of the structural component can be estimated using material properties and local stress information. A key input in this step is the so-called S–N curve, which describes the relationship between the stress (typically the stress amplitude) and the number of cycles to failure for a given material and stress concentration factor, $k_{t}$. The factor $k_{t}$ quantifies how much the local stress is amplified due to defects or geometric features such as holes, fillets, or notches. Based on these parameters, along with the load history, the mission definition and the principal structural element (PSE), the fatigue life, quantified as number-of-flights-to-failure NN, can be formally expressed as:
$N=\mathscr{F}(\text{material},k_{t},\text{historical data},\text{PSE},\text{mission data},\text{stress data}).$ (1)
A linear damage accumulation model (Miner’s rule [bland1946discussion]) is finally adopted to estimate fatigue damage. The core idea is that the damage caused by each stress cycle is independent and additive, and thus the total damage is simply the sum of these contributions. Under this model, the damage $\mathbb{D}$ accumulated after $n$ flights per mission is simply given by:
$\mathbb{D}=\frac{n}{N},$ (2)
such that by definition, failure occurs when the number of flights reaches the critical value $n^{*}=N$, i.e. at $\mathbb{D}=1$. Although this model simplifies the effects of plasticity and the interactions between loads, it remains widely used in aerospace applications and serves as the reference for ground truth values in this study.
2.2 Fatigue life prediction using Machine Learning
As previously outlined, fatigue life estimation relies on a sequential process that demands substantial computational and human resources. It typically requires a combination of high-fidelity CFD and FEM simulations, followed by load cycle reconstruction and counting techniques. This workflow is not only time-consuming but also difficult to scale. In contrast, machine learning offers the potential for near-instantaneous inference and significant reductions in both computational cost and operational complexity, hence the opportunity.
The rationale starts from assuming that $\mathscr{F}(\cdot)$ in Eq. 1 is a deterministic function, and aim to approximate it using a suite of ML models. Now, after preliminary inspection it was clear that modeling $\mathscr{F}$ with a single neural network led to stagnant loss curves and suboptimal learning. This, together with the fact that including every step of the process into a single neural network hinders interpretability, led us to reformulate the solution as a sequential pipeline composed of three supervised learning (regression) tasks, such that the global task (approximate $\mathscr{F}$ in Eq. 1) can be seen as a function composition, i.e. conceptually we have $\mathscr{F}=f_{3}\circ f_{2}\circ f_{1}$, where $f_{1}$, $f_{2}$ and $f_{3}$ are nonlinear functions mapping different stages of the fatigue damage assessment pipeline described in Fig. 1. As we shall see, each of these functions are approximated by different ML and statistical models.
As in the traditional approach, the ML pipeline begins by translating the mission into a structured sequence of flight segments, each described by 12 constant flight parameters. These parameters capture the operational characteristics of the aircraft during each segment (e.g., taxi, climb, cruise) and serve as input features for the first stage of the pipeline. For each $\{\text{mission},\text{PSE}\}$ pair, the pipeline proceeds in three phases:
  • 1.
    Phase I – Stress prediction: The objective of this phase is to predict the stress vector $\mathbf{S}\in\mathbb{R}^{4}$ at each principal structural element (PSE) and for every mission segment, using the corresponding flight parameters $\mathbf{P}\in\mathbb{R}^{12}$ as input. Details on the components of $\mathbf{P}$ and $\mathbf{S}$ are provided in A.1.
    Formally, this task would involve learning a function $f_{1}$ such that
    $f_{1}:\mathbb{R}^{12}\rightarrow\mathbb{R}^{4},\quad\mathbf{S}=f_{1}(\mathbf{P},\text{PSE})$
[…]
  • 2.
    Phase II – Damage prediction: The second phase of the pipeline estimates, from the stress data, the fatigue damage accumulated at each PSE over $n$ flights of a given mission. Since the particularities of each PSE are already encoded in the different stresses, PSE is not anymore an input variable at the stage. Now, there are two distinct damage components: $\mathbb{D}_{\text{GAG}}$, the total (accumulated) damage caused by the Ground–Air–Ground (GAG) cycle after $n$ flights, and $\mathbb{D}_{\text{G\&M}}$, the total (accumulated) damage induced by gusts and maneuver loads during flight segments after $n$ flights. These are the two primary contributors to fatigue degradation in metallic aircraft structures. Since they are distinct, we separate the analysis and consider two separate tasks:
    (i) The first task is that of predicting the typical ‘per-flight’ damage caused by the Ground–Air–Ground (GAG) cycle.Defining $\mathscr{D}_{\text{GAG}}$ as a latent random variable that characterises the GAG damage within a single flight, our first task is predicting the average value of this latent random variable, $\overline{\mathscr{D}_{\text{GAG}}}$, such that later we will use $\mathbb{D}_{\text{GAG}}=\overline{\mathscr{D}_{\text{GAG}}}\cdot n$ to estimate the total GAG damage (i.e. $\mathbb{D}_{\text{GAG}}$ is a random variable defined as the sum of $n$ independent and identically distributed random variables $\mathscr{D}_{\text{GAG}}$). The prediction of $\overline{\mathscr{D}_{\text{GAG}}}$ uses as input variables the average stress values during ground and flight segments $\bar{\mathbf{S}}_{\text{ground}}\in\mathbb{R}$ and $\bar{\mathbf{S}}_{\text{flight}}\in\mathbb{R}^{4}$ (computed from the Phase I stress predictions), along the time the aircraft spends in flight and on the ground during a single flight, $t_{\text{ground}}$, $t_{\text{flight}}\in\mathbb{R}$ and the stress concentration factor, $k_{t}\in\mathbb{R}$ (see B for details). In principle this is a per-flight quantity and thus independent of $n$. However, although $n$ is large ($n\in[800,13{,}000]$) and the average damage is expected to asymptotically converge to the true per-flight value-as per the law of large numbers, see C for more details-we also add $n$ as an input variable, to account to possible deviations from the true mean due to finite samples effects. Mathematically
    $f_{2}^{\text{GAG}}:\mathbb{R}^{9}\rightarrow\mathbb{R},\quad\overline{\mathscr{D}_{\text{GAG}}}=f_{2}^{\text{GAG}}(\bar{\mathbf{S}}_{\text{ground}},\bar{\mathbf{S}}_{\text{flight}},t_{\text{flight}},t_{\text{ground}},k_{t},n).$
    As previously stated, the total GAG-related damage after $n$ flights is $\mathbb{D}_{\text{GAG}}=\overline{\mathscr{D}_{\text{GAG}}}\cdot n$.
    (ii) Likewise, the second task is that of predicting the the typical, per-flight damage caused by the Gusts and Maneuvers (G&M) cycles $\overline{\mathscr{D}_{\text{G\&M}}}$. This prediction uses the same inputs as before, except for the ground-related variables. Mathematically, we have
    $f_{2}^{\text{G\&M}}:\mathbb{R}^{7}\rightarrow\mathbb{R},\quad\overline{\mathscr{D}_{\text{G\&M}}}=f_{2}^{\text{G\&M}}(\bar{\mathbf{S}}_{\text{flight}},t_{\text{flight}},k_{t},n).$
    The total G&M-related damage after $n$ flights $\mathbb{D}_{\text{G\&M}}=\overline{\mathscr{D}_{\text{G\&M}}}\cdot n$.
    We train two different MLPs to approximate $f_{2}^{\text{GAG}}$ and $f_{2}^{\text{G\&M}}$. Note that the use of averaged stress values instead of segment-wise cycle estimation is motivated by the high variability of load spectra across flights and the lack of deterministic predictability of individual stress cycles. Since missions typically involve a large number of flights ($n\in[800,13{,}000]$), this formulation is justified by the Central Limit Theorem and the assumption of independent (Miner’s rule), identically distributed flights (see C for more details). Both models were trained using log-transformed damage values to address the wide dynamic range in the output space, and we use a Mean Absolute Error (MAE) loss function to penalize poor performance on high-damage cases, which are the most critical in fatigue life estimation. Full details on the data preparation, feature engineering, model architecture and split strategy are provided in B.
  • 3.
    Phase III - Life prediction: The final phase computes the fatigue life NN defined as the number-of-flights-to-failure for a given mission and PSE. The model takes as inputs the number of flights, $n$, and the two sources of accumulated fatigue damage $\mathbb{D}_{\text{GAG},\mathbb{D}_{\text{G\&M}}}$ predicted in Phase II, such that $N=f_{3}(\mathbb{D}_{\text{GAG}},\mathbb{D}_{\text{G\&M}})$. Rather than approximating $f_{3}$ via an additional ML model, we instead apply the simpler Miner’s rule directly, as described in subsection 2.1. This linear damage accumulation model assumes failure occurs when total damage reaches unity. Accordingly, the total damage $\mathbb{D}=\mathbb{D}_{\text{GAG}}+\mathbb{D}_{\text{G\&M}}$ and thus fatigue life is computed as:
    $N=f_{3}(n,\mathbb{D}_{\text{GAG},\mathbb{D}_{\text{G\&M}}})=\frac{n}{\mathbb{D}_{\text{GAG}}+\mathbb{D}_{\text{G\&M}}}.$ (3)
2.3 Dataset and split description
Our dataset contains information regarding 38 wing PSEs, labeled 1, 2, $\dots$ to 38, four different $k_{t}$ values, and seven different missions operated by a specific certification of the studied tactical aircraft, labeled A, B, $\dots$ to G. Each mission represents accumulated data over a large number of flights, typically ranging between $n\in[800,13000]$.
As stated in subsection 2.1, fatigue damage assessment requires conducting different steps (Figure 1), such as running FEM simulations, load cycles construction and finally, computing damage (and life) with the S-N curves. Our database, contains the output of each one of these steps, which will serve as the ground truth values for our ML approach.
As previously discussed, missions are divided into operational segments such as taxi, climb, cruise, descent, and approach, being each mission segment described in the dataset by 12 scalar flight parameters collected into a vector $\mathbf{P}\in\mathbb{R}^{12}$. These flight parameters induce certain stresses at each PSE (FEM output), represented by 4-dimensional stress vectors $\mathbf{S}\in\mathbb{R}^{4}$, which represent the local stress at the component for different loading conditions. Details about P and S provided in A.1. We leverage this data to address phase I from subsection 2.1.
In addition, for all the missions, the load cycles induced by gusts and maneuvers during every segment, derived from statistical load distributions (subsection 2.1), are available at each PSE, along with the GAG cycles and the respective fatigue damages (for the four $k_{t}$ values) accumulated over $n$ flights of the respective mission. The ground truth of fatigue life is easily computed with these damages.
Data split – Instead of performing a random split on all the $38\times 7\times 4=1064$ $\{\text{PSE},\text{Mission},k_{t}\}$ combinations- which could result in data leakage and in some PSEs being absent from the test set -the following split was performed. For each PSE, a set of five out of the seven missions have been randomly selected for the training set (for the four $k_{t}$ values, treated as numerical variables), one mission for the validation set and another mission for the test set. The validation set is used for hyperparameter tuning while the test set is used to assess the overall performance of the pipeline. Details on how this split is applied in each phase are provided in A.2 and B.3.
2.4 Note on software
ML models were developed using pyLOM [pyLOM], a high-performance computing-enabled framework for parallel reduced-order modeling implemented in PyTorch [paszke2019pytorch]. In addition, we also used standard Python libraries, such as scikit-learn [scikit-learn] for polynomial regression and data preprocessing, pandas [mckinney-proc-scipy-2010] and NumPy [harris2020array] for dataset manipulation, and matplotlib [Hunter:2007] for visualization.
3 Results
In this section, we start by presenting the results of every step of the pipeline $f_{1},f_{2},f_{3}$, along with a thorough error quantification of each of the predictions. The relative error used throughout this section is the absolute percentage error:
$\text{Relative Error (\%)}=\left|\frac{\hat{y}-y}{y}\right|\times 100,$ (4)
where $\hat{y}$ and $y$ are the predicted and true values respectively.
3.1 Phase I: Stress predictions
First, we will study the results of the stress vector S predictions of every mission segment and for every PSE $\in$ Test set. This phase approximates $f_{1}$ in Fig. 1.
Mean Relative Error (%) Std Q1 Q2 Q3
$1g$ 0.04 0.03 0.019 0.03 0.06
Table 1: Relative Error statistics for S predictions for ground segments, displaying the mean relative error (relative error averaged over all PSEs and ground segments), the standard deviation, and the first three quartiles.
[…]
where $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$ are the polynomial coefficients, obtained by minimizing the residual sum of squares between predictions and ground truth values (ordinary least squares) [scikit-learn].
First, in Table 1 we display the statistics of the relative error for the $1g$ stress prediction in the test set. The mean relative error (MRE) –i.e. the relative error averaged over all ground segments and PSEs– is substantially below $1\%$, indicating a very good performance. To gain insight on the dependency of the prediction error on the PSEs and missions, we depart from pointwise statistics and in Fig. 3 we plot the Mean Relative Error (MRE) of the $1g$ stress predictions conditioned on (i.e. grouped by) PSEs. As observed, although there is a slight variability in performance between PSEs, for all PSEs the MRE systematically remains below 0.1%. Similarly, Table 2 shows the performance (again in terms of MRE) but now grouped by missions instead of PSEs. As observed, all the missions exhibit excellent performance and again, MRE systematically remains below 0.1%. Overall, these results indicate that the dependence of the $1g$ stresses on flight parameters for ground segments is very simple and is well captured by a second order polynomial.
Figure 3: Mean relative error (in $\%$) for the predicted $1g$ for ground segments, grouped by PSEs.
A B C D E F G
$1g$ 0.05 0.05 0.03 0.02 0.04 0.04 0.05
Table 2: MRE for the predictions of each element of the stress vector S over ground segments, grouped per mission class.
Figure 4: Learning curves (training and validation loss) for the MLP that approximates $f_{1}^{\text{flight}}$.
[…]
MRE (%) Std Q1 Q2 Q3
$1g$ 0.86 0.89 0.29 0.63 1.06
$\Delta\text{vman}$ 0.48 0.50 0.14 0.33 0.68
$\Delta\text{vgust}$ 0.73 0.66 0.25 0.56 1.03
$\Delta\text{turn}$ 1.17 1.05 0.40 0.87 1.60
Table 3: Relative error statistics for the predictions of all four elements of the stress vector S over flight segments: mean relative error (MRE, %, averaging over all PSEs and flights segments), standard deviation, and the first three quartiles Q1 (percentile 25), Q2 (median) and Q3 (percentile 75).
(a) Mean relative error for $1g$ predictions for flight segments grouped by PSEs
(b) Mean relative error for $\Delta\text{vman}$ predictions for flight segments grouped by PSEs
(c) Mean relative error for $\Delta\text{vgust}$ predictions for flight segments grouped by PSEs
(d) Mean relative error for $\Delta\text{turn}$ predictions for flight segments grouped by PSEs
Figure 5: Mean relative error grouped by PSEs for the all the predicted stresses
Mission A B C D E F G
$1g$ 1.12 0.60 0.79 0.66 0.65 0.65 0.66
$\Delta\text{vman}$ 0.68 0.37 0.36 0.42 0.41 0.42 0.30
$\Delta\text{vgust}$ 0.90 0.52 0.68 0.70 0.53 0.69 0.57
$\Delta\text{turn}$ 1.29 0.56 1.66 0.97 0.48 0.64 0.75
Table 4: MRE per mission for S predictions over flight segments
Figure 6: Histograms of input variables for mission A vs all the other missions. Specific support ranges have been hidden for confidentiality reasons.
[…]
3.2 Phase II: GAG and G&M Damage prediction
In this phase, both sources of fatigue damage are predicted separately with two different MLP models, that approximate $f_{2}^{\text{GAG}}$ and $f_{2}^{\text{G\&M}}$ respectively. In Fig. 7 the learning curves of the models are depicted. They were trained for 5000 epochs (note that working with mini-batches results in many iterations per epoch). As observed, during the training process, the validation loss remains close to the training loss (MAE) for both models. The gusts and maneuvers damage model shows a more noisy learning process than the GAG damage model, but still manages to learn. Finally, the validation sets remain flat, so extending the training for more epochs might lead to a high-variance scenario, where the model would learn spurious noise.
(a) GAG model learning curves
(b) Gusts and maneuvers model learning curves
Figure 7: Learning curves for both damage models approximating (a) $f_{2}^{\text{GAG}}$ and (b) $f_{2}^{\text{G\&M}}$.
In Figure 9 and Figure 9 we show a summary of the relative error (%) statistics of the damage predictions $\mathbb{D}_{\text{GAG}}$ and $\mathbb{D}_{\text{G\&M}}$, for all the points in the test set. As we can see, the mean relative error is moderately large and substantially larger that the median, what suggests a skewed, fat-tailed error distribution. Notice, indeed, that the third quartile is not symmetric with respect to the first quartile, and that the max error is disproportionally large. This suggests that pointwise error statistics are being deformed –and the average performance apparently decreased– due to the presence of outliers. So it is important to ascertain where these outliers are located. Since the models predict fatigue damage –which can lead to instantaneous breakage–, greater damages represent more critical situations in both cases, as fatigue failure will be achieved earlier, resulting in shorter fatigue lives. Therefore, it is in the high-damage region that the MLP models must be more accurate, and therefore we need to certify that outliers are never present in this region.
Relative error stats Values
Mean 8.12
Standard deviation 12.10
First quartile (25%) 1.21
Median (50%) 3.67
Third quartile (75%) 9.40
Min error 0.08
Max error 80.95
Figure 8: Relative error statistics for the prediction of $\mathbb{D}_{\text{GAG}}$.
Relative error stats Values
Mean 9.49
Standard deviation 21.80
First quartile (25%) 2.47
Median (50%) 4.94
Third quartile (75%) 8.40
Min error 0.03
Max error 202.29
Figure 9: Relative error statistics for the prediction of $\mathbb{D}_{\text{G\&M}}$.
Figure 10: Scatter plot of the relative error of $\mathbb{D}_{\text{GAG}}$ as a function of its ground truth value, for every point in the test set.
Figure 11: Scatter plot of the relative error of $\mathbb{D}_{\text{G\&M}}$ as a function of its ground truth value, for every point in the test set.
Now, recall that we are using the mean absolute error (MAE) as the loss function in both cases
$\mathscr{L}_{\text{MAE}}=\frac{1}{m}\sum_{i=1}^{m}\left|y_{i}-\hat{y}_{i}\right|,$ (5)
where $m$ are the number of samples on the mini batch and $y_{i}$, $\hat{y}_{i}$ are the true and predicted damages, respectively. By construction, this loss function already penalizes error for higher damages than for lower damages. To confirm this dependence of the prediction’s error on the actual damage, in Figure 10 and Figure 11 we plot, in log-log scales, the relative error for the prediction of accumulated damages $\mathbb{D}_{\text{GAG}}$ and $\mathbb{D}_{\text{G\&M}}$ as a function of their ground truth values. Each point in these two scatter plots represents a $\{\text{Mission},\text{PSE},k_{t}\}\in\text{Test set}$. Clearly, model’s performance is better for higher damages (more critical situations), and it is in the region of low damages that the error outliers tend to concentrate (observe that the plots are logarithmic). This bias is even stronger for the GAG damage predictions –which is actually the main source of fatigue damage of the studied aircraft–, showing a statistically significant weak inverse correlation between the accumulated damage and the relative error of the prediction (Spearman’s $\rho=-0.353$ with p-value $6.7\times 10^{-5}$). In contrast, for the accumulated damage caused by gusts and maneuvers, the correlation is very weak and not statistically significant (Spearman’s $\rho=-0.17$ with p-value =0.085=0.085). However, the distribution outliers (e.g., following Tukey fences, those points with relative error larger than $Q3+1.5IQR=17.3\%$) are mainly located in the low-damage range.
3.3 Phase III: Fatigue life results
Miner’s Rule is finally applied to obtain total fatigue damage after the $n$ flights, and with Eq.3 fatigue life NN can easily be calculated. Since the MLP models for $f_{2}^{\text{GAG}}$ and $f_{2}^{\text{G\&M}}$ showed poorer performance for lower fatigue-related damages, errors propagate and thus we expect to get less accurate predictions of fatigue life for higher fatigue lives, since $N\propto\mathbb{D}^{-1}$.
In Figure 12 we scatter plot the relative error of each fatigue life prediction NN as a function of its ground truth values. As observed, errors of the predictions are not uniformly distributed along fatigue life. For lives over $10^{6}$ the model shows poorer performance. This behavior is similar to that of the inverse of Figure 10. This is expected as the main source of damage for the studied aircraft lies in the GAG cycle. This poorer performance for lives over $10^{6}$ flights does not represent a significant limitation for our model, since these very long lives do not represent critical situations and big errors do not pose a safety hazard. In contrast, shorter lives are more critical, as failure will be reached earlier in the aircraft operation.
Figure 12: Scatter plot of the relative error of fatigue life prediction as a function of its ground truth value, for every point in the test set.
For completeness, in Figure 13 we also show the mean relative error of fatigue life prediction conditioned on $k_{t}$, mission and PSE, respectively. As observed in 13(a), cases with smaller $k_{t}$ concentrate larger relative error. This is a sensible result: small $k_{t}$ values in general tend to have longer fatigue lives (i.e. the region where the ML model performs worse) because in these cases the local maximum stress is barely amplified by defects or geometry characteristics.
Likewise, in 13(c) we observe that some PSEs show poorer fatigue life prediction, but this is simply because these are subjected to lower stresses, leading to higher fatigue lives and thus poorer model’s performance. Incidentally, note that PSE 15 was not included in the plot as its high error was off-scale. Finally, 13(b) shows only slight variability in performance across missions.
(a) MRE of fatigue life prediction NN, for all test set data grouped by their $k_{t}$ values.
(b) MRE of fatigue life prediction NN, for all test set data grouped by their mission class.
(c) MRE of fatigue life prediction NN, for all test set data grouped by PSEs.
Figure 13: Mean relative error for the fatigue life prediction NN, grouped by (a) $k_{t}$, (b) mission class, and (c) PSEs.
Finally, Table 6 depicts a summary of the main relative error statistics for fatigue life predictions and in Table 6, the same statistics are presented excluding samples with lives exceeding $10^{6}$ flights, since these are non-critical and some of them can highly perturb the error statistics. In this latter region, performance is highly improved and predictions become more accurate.
Table 5: Full domain
Relative error stats Values
Mean 11.37
Standard deviation 39.48
First quartile (25%) 1.05
Median (50%) 3.60
Third quartile (75%) 9.02
Min error 0.07
Max error 424.89
Table 6: $N\in(10^{3},10^{6})$
Relative error stats Values
Mean 3.99
Standard deviation 4.03
First quartile (25%) 0.88
Median (50%) 2.88
Third quartile (75%) 5.34
Min error 0.067
Max error 17.65
Figure 14: Prediction interval for the relative error
3.4 Uncertainty quantification: prediction intervals
Since our ML pipeline aims to provide a solution for industrial and safety-critical applications, authorities require not only accuracy, but confidence in its predictions. In fatigue analysis, it us customary to use scatter factors to account for the variability in material fatigue properties and the uncertainties inherent in predicting fatigue life. To that end, following the principles of [lacasa2025certification] we now build prediction intervals for fatigue life predictions.
We start by considering the relative error of the predictions of the calibration set in the region $N\in(10^{3}-10^{6})$ to avoid impractical situations that can perturb the error statistics. Since the relative error is larger or equal to zero, we construct bootstrapped confidence interval of 95th percentile (CI95) of this metric [lacasa2025certification]. In Figure 14 we plot the bootstrap average of this percentile, such that the range $0,\overline{P95}]$ in principle covers roughly the relative error of about $95\%$ of the points in the calibration and test set. The analysis showed that, using the validation set described in subsection 2.3 as the calibration set to build a prediction interval for the 95% coverage, we achieve a coverage in the test set of 95.24%, CI95 [88.25, 98.69]. This corresponds to an upper bound of the relative error of 13.50%.
Consequently, when predicting the life of a new, previously unseen sample, the true fatigue life can be assured to be $N\in[\hat{N}(1-\varepsilon),\hat{N}(1+\varepsilon)]$ — where $\varepsilon=0.135$, and $\hat{N}$ is the predicted fatigue life — with approximately 95% confidence.
3.5 A posteriori training/test split analysis
Before assessing the performance of the different ML models, it is a good practice to check whether the training/test data split is adequate, as incorrect splits can be a source of p-hacking –which misleadingly amplifies the performance on the test set– or poor generalization [lacasa2025certification]. Now, since in this work the output data of model $f_{1}$ is the input data of model $f_{2}$, we now perform such data split analysis a posteriori, and focus in the input data of MLP models $f_{2}$ and $f_{3}$, as the configuration space is smaller than the one for $f_{1}$ (note that this analysis was performed for all steps):
$N=f_{3}(n,\overline{\mathscr{D}_{\text{GAG}}},\overline{\mathscr{D}_{\text{G\&M}}})=f_{3}(n,f_{2}(\bar{\textbf{S}}_{flight},\bar{\textbf{S}}_{ground},t_{flight},t_{ground},k_{t},n)),$
where $f_{2}=\{f_{2}^{\text{GAG}},f_{2}^{\text{G\&M}}\}$. Following [lacasa2025certification], we consider a split to be adequate if (i) finding the same data samples appearing in both the train and test sets has vanishingly small probability, (ii) the feature vector distributions of both sets are similar, and (iii) the test set data is adequately close from the training set.
Criterion (ii) is checked by performing 2-sample hypothesis tests (2-sample Kolmogorov-Smirnov (KS) [scholz1987k], 2-sample Anderson-Darling (AD) [hodges1958significance] and 2-sample chi-2 tests) on marginal training and test distributions for each input variable. Criterion (iii) is implemented in the so-called tesselation and proximity method [lacasa2025certification], by detecting test set points that are either excessively close (statistically speaking) to the training set (source of p-hacking), or statistically too far (isolated points with poor generalization).
Figure 15: Training vs test marginal distributions for each numerical input variable.
Figure 16: Training vs test marginal distributions for each categorical input variable.
Figure 17: Training vs test marginal distributions for the output variable NN.
In Figure 15 we plot the marginal training and test set distributions for each numerical input variable, whereas in Figure 16 we plot the ones for the categorical variables. Although variables in Figure 16 are actually numerical, for the sake of this analysis they are treated as categorical as they only take a small set of possible values. Therefore, they cannot be considered continuous and a different test is applied ($\chi^{2}$). Likewise, Figure 17 presents the marginal training and test distributions for the output variable NN. Visually, the training and test distributions are similar, although small differences exist. To quantify the significance of these, we have run several hypothesis tests, where the null hypothesis H0 is that, for each variable, each pair of populations are drawn from the same distribution. Since each hypothesis test has its own null hypothesis, no correction for multiple testing is required, and we adopt the standard choice for rejecting the null. In Table 7 we present the p-values of these tests, where the null hypothesis cannot be rejected for any variable except for $n$ and $t_{flight}$. Note nonetheless that, as observed in Figure 16, the range values for these two variables is similar in both sets, indicating that at least there are no isolated values in $n$ or $t_{flight}$ in the test set.
Besides, from the 121 samples in the test set, the proximity method flagged three points (2.48% of the test set) as p-hacking and eight as isolated samples (6.6% of the test set). As expected, the p-hacking points showed overly optimistic predictions, with a mean relative error of 0.76%, whereas this value for the isolated points is 10.21%.
Finally, we can conclude that the test set is a good representation of the training set in both the input and output spaces and that the split is adequate, hence providing further assurance that the error and uncertainty quantification developed in previous sections represents well the performance of the model.
Test $k_{t}$ $\bar{1g}_{flight}$ $\bar{1g}_{ground}$ $\bar{\Delta_{Vgust}}$ $\bar{\Delta_{Vman}}$ $\bar{\Delta_{Turn}}$ $n$ $t_{flight}$ $t_{ground}$ NN
KS - 0.49 0.25 0.08 0.89 0.38 - - - 0.95
AD - 0.25 0.25 0.25 0.25 0.25 - - - 0.25
$\chi$² 1.00 - - - - - 0.001 0.001 0.92 -
Table 7: p-values for the two-sample Kolmogorov-Smirnov, Anderson-Darling and $\chi^{2}$ tests, comparing the empirical distribution of each input and output variables in the training vs test sets. Results indicate that, except for $n$ and $t_{\text{flght}}$, for the rest of input variables the null hypothesis of equal distribution cannot be rejected.
4 Conclusions
This study presents a modular machine learning-based pipeline for fatigue life prediction at different wing locations, across multiple stress concentration factors ($k_{t}$), using mission-level flight parameters as input. The framework offers a fast, scalable, and accurate complement to traditional simulation-based approaches, with direct applications in early-stage aircraft design, mission planning, and maintenance strategies. The pipeline combines expert-domain feature engineering with deep learning models tailored to flight and ground segments. Once trained, it enables instant fatigue life and damage predictions which turn to be particularly accurate in the medium-to-low fatigue life regime, which is the critical regime for design decisions. The use of predicted stress components as intermediate variables supports both accuracy and physical interpretability. The pipeline has been statistically validated, making it consistent with ongoing initiatives to create certifiable ML tools in safety-critical aerospace systems.
Looking ahead, while the current study is limited to wing components due to data availability, the methodology is extensible to other materials and aircraft structures, such as the fuselage and stabilizers. Future work should also focus on broader structural coverage, improved stress modeling, modifications to the pipeline aimed at promoting conservativeness, (e.g., using modified loss functions to penalize non-conservative predictions), and further integration of uncertainty quantification conditioned on input space [lacasa2025certification].
In summary, the proposed approach represents a novel and effective contribution to fatigue life modeling, demonstrating the feasibility of using machine learning to complement and accelerate conventional analysis pipelines in aerospace structural integrity.
Acknowledgments – The authors acknowledge funding from project TIFON (PLEC2023-010251) funded by MCIN/AEI/10.13039/501100011033, Spain. $L$ acknowledges partial support from projects MISLAND (PID2020-114324GB-C22), project CSxAI (PID2024-157526NB-I00) funded by MICIU/AEI/10.13039/501100011033/FEDER, UE, project MdM Seal of Excellence (CEX2021-001164-M) funded by MICIU/AEI/10.13039/501100011033 and from the European Commission Chips Joint Undertaking project No. 101194363 (NEHIL). Finally, all authors gratefully acknowledge the Universidad Politécnica de Madrid (www.upm.es) for providing computing resources on Magerit Supercomputer.
Appendix A Additional details for Phase I
A.1 Flight parameters and stress vectors
Each segment is parametrized by 12 scalar flight parameters, collected in a vector $\mathbf{P}\in\mathbb{R}^{12}$ whose element description is detailed in Table 8. We will use these parameters to raise the input of this phase. On the other hand, for each of the 38 wing locations (PSEs), stress variables are collected in a four-dimensional stress vector $\textbf{S}=(1g,\ \Delta{\text{vman}},\ \Delta{\text{vgust}},\ \Delta{\text{turn}})$, which represents the local stress at the component for different loading conditions. The physical meaning of each variable in $\mathbf{S}$ is detailed in Table 9. Note that for ground segments $\Delta{\text{vman}}=\ \Delta{\text{vgust}}=\Delta{\text{turn}}=0$.
Parameter Description
Flaps Degree of extension of the flaps in the segment.
TAS True airspeed during the segment.
Altitude Altitude at which the segment takes place.
Time Time spent in the segment.
Distance Distance covered in the segment.
Thrust Engine thrust during the segment.
Pressure Pressure difference at the segment altitude.
Mass Aircraft mass during the segment.
%CMA Relative position of the aircraft center of gravity with respect to the mean aerodynamic chord.
ZFW Zero fuel weight of the aircraft.
PL Payload weight.
FW Fuel weight.
Table 8: Description of the 12 scalar flight parameters $\mathbf{P}\in\mathbb{R}^{12}$ for each segment
Stress Component Description
$1g$ Equilibrium stress at the studied PSE corresponding to steady flight.
$\Delta_{\text{Vman}}$ Incremental stress caused by a 1.5g vertical maneuver at the studied PSE.
$\Delta_{\text{Vgust}}$ Incremental stress caused by a 1.5g vertical gust at the studied PSE.
$\Delta_{\text{Turn}}$ Incremental stress caused by a 1.5g turn at the studied PSE.
Table 9: Description of the four stress components $\mathbf{S}\in\mathbb{R}^{4}$ for each segment
A.2 Split
As stated in subsection 2.3, for each PSE, five missions were selected for training, one for validation and one for testing. However, in this phase, each {segment, PSE} constitutes a sample. This implies that, for each PSE, all the segments belonging to the five training missions are samples in the training set, and all the segments belonging to the validation and test missions belong to the validation and test sets respectively. This yields the split presented in Table 10
Train size Validation size Test size
Flight segments 1644 310 344
Ground segments 380 76 76
Table 10: Set sizes for Phase I.
A.3 Input, output variables and model architecture
[…]
For flight segments (i.e. climb, cruise, descent, approach), wings are also subjected to aerodynamic loads, so more variables will impact on stresses (for instance, lift is proportional to TAS2). So the following set of features are used as input variables for $f_{1}^{\text{flight}}$: {PSE (binary encoded), Flaps, Altitude, TAS, Mass, FW, Thrust}. All of them were normalized with a regular MinMax Scaler.
[…]
Finally, the hyperparameters of the deep neural network used to approximate $f_{1}^{\text{flight}}$ and other optimization parameters are depicted in Table 11. Hyperparameter tuning was performed manually on the validation set to ensure generalization.
Train/Val/Test size 1664/310/344
Hidden layer sizes 50/50
Activation Function ReLU
Loss Function MAE
Optimizer Adam
Weights initialization Xavier
Learning rate $\alpha$ $8\cdot 10^{-3}$
Epochs 1000
Regularization None
Mini batch Size 256
Scheduler: $\gamma$, Step 0.975, 30
Table 11: Hyperparameters and optimization choices for the deep MLP used to approximate the stress properties for any given PSE for flight segments ($f_{1}^{\text{flight}}$).
Appendix B Phase II: Additional details for Phase II
B.1 Data preprocessing
Predicted stresses from all the mission segments are averaged to form
$\bar{\mathbf{S}}_{\text{ground}}=\langle 1g\rangle\in\mathbb{R}$
for ground segments and
$\bar{\mathbf{S}}_{\text{flight}}=(\langle 1g\rangle_{\text{flight}},\langle\Delta\text{vman}\rangle_{\text{flight}},\langle\Delta\text{vgust}\rangle_{\text{flight}},\langle\Delta\text{turn}\rangle_{\text{flight}})\in\mathbb{R}^{4}$
for flight segments. In addition, the durations associated with each mission segment are used to compute $t_{flight}$ and $t_{ground}$. All the averages $\langle\cdot\rangle$ above are computed in a similar way: as an example, the averaged equilibrium stress during flight segments
$\langle 1g\rangle_{flight}=\frac{1}{t_{flight}}\sum 1g_{i}\cdot t_{i},$
where $1g_{i}$ and $t_{i}$ are the predicted equilibrium stress of flight segment $i$ and its duration, respectively.
B.2 Database
Load cycles: For each segment, load cycles caused by vertical maneuvers, gusts, and turns are included. In addition, the $n$ GAG cycles corresponding to the $n$ flights are also included. As detailed in subsection 2.1, these are stored as a vector with three components: Maximum stress, Minimum stress and Number of occurrences (after $n$ flights). The associated damage for each of these cycles is also available (for the four $k_{t}$ values).
B.3 Split
As stated in subsection 2.3, for each PSE, five missions were selected for training, one for validation and one for testing (with their respective four $k_{t}$ values). In this phase, every {mission, PSE, $k_{t}$} represents a sample. After filtering extreme values and unrealistic data, the set sizes for the GAG model and the gusts and maneuvers model are presented in Table 12.
Table 12: Set sizes
Train size Validation size Test size
GAG 621 125 123
Gusts and maneuvers 533 104 107
B.4 $f_{2}^{\text{GAG}}$: input, output variables and model specs
$f_{2}^{\text{GAG}}$ takes 9 numerical input variables: $k_{t},\bar{\mathbf{S}}_{\text{flight}}\in\mathbb{R}^{4},\bar{\mathbf{S}}_{\text{ground}},n,t_{\text{flight}},t_{\text{ground}}$. Note that as only wing PSEs are studied and all the wing PSEs have the same material, the material is not an input variable. The physical rationale behind each one of these input variables is as follows:
  • 1.
    $k_{t}$: The stress concentration factor, $k_{t}$, has a strong detrimental effect on fatigue life. This variable quantifies the amplification of local stress caused by defects or geometric characteristics, such as holes or joints. High $k_{t}$ values promote crack nucleation. It is defined as:
    $k_{t}=\frac{\text{Local maximum stress}}{\text{Nominal stress}}$ (6)
    and it is a key variable in this problem.
  • 2.
    $\langle 1g\rangle_{\text{ground}}$ and $\langle 1g\rangle_{\text{flight}}$: about the form and amplitude of the GAG cycle.
  • 3.
    $\langle{\Delta\text{vgust}}\rangle,\langle{\Delta\text{vman}}\rangle,\langle{\Delta\text{turn}}\rangle$: these averaged incremental stresses provide information about the magnitude of the stress oscillations caused by gusts and maneuvers. In the GAG cycle, the maximum stress will always be greater than $\langle 1g\rangle_{\text{flight}}$. The same happens with the minimum stress.
  • 4.
    $n$, $t_{flight}$, $t_{ground}$: These variables have a statistical impact. Longer flights increase the likelihood of encountering greater gusts or executing more demanding maneuvers, which augments the amplitude of the GAG cycle and, consequently the average GAG damage, $\overline{\mathscr{D}_{\text{GAG}}}$, which is the output of the model. Besides, since the number of samples is finite, $n$ is included to account for potential deviations from the true mean damage.
Input and output variables were normalized with MinMax Scaler and a logarithm was applied to the damage due to the scale difference between samples.
Finally, the specs of the MLP model and its hyperparameters are shown in Table 13. Hyperparameter tuning was performed manually with the validation set to ensure generalization.
Train/Validation/Test sizes 621/125/123
Hidden layer sizes 64/64
Activation Function tanh
Loss Function MAE
Optimizer Adam
Weights initialization Xavier
Learning rate $\alpha$ $8\cdot 10^{-3}$
Epochs 5000
Regularization Dropout, 0
Mini batch Size 128
Scheduler: $\gamma$, Step 0.975, 30
Table 13: Specifications and hyperparameters of the MLP of $f_{2}^{\text{GAG}}$.
B.5 $f_{2}^{\text{G\&M}}$: input, output variables and model specs
In this case we only have 7 input variables: $k_{t},\bar{\mathbf{S}}_{\text{flight}}\in\mathbb{R}^{4},n,t_{flight}$. The rationale of each variable is similar to that for $f_{2}^{\text{GAG}}$; $\langle 1g\rangle_{\text{flight}}$ shows the stress level around which the oscillations that cause this damage take place. The magnitude of these oscillations is captured by $\langle{\Delta\text{vgust}}\rangle,\langle{\Delta\text{vman}}\rangle,\langle{\Delta\text{turn}}\rangle$ and $n$, $t_{flight}$ have the same statistical impact. Besides, longer flights typically imply more gusts encountered and maneuvers executed (not necessarily larger) and consequently greater damage. The output variable is the per-flight damage $\overline{\mathscr{D}_{\text{G\&M}}}$, such that the cumulated damage is simply $\mathbb{D}_{\text{G\&M}}=n\cdot\overline{\mathscr{D}_{\text{G\&M}}}$. Once again, variables were normalized with MinMax Scaler and logarithm was applied to the damage. The MLP specs and hyperparameters are specified in Table 14. Hyperparameter tuning was performed manually with the validation set to ensure generalization.
Table 14: MLP architecture
Train/Validation/Test size 533/104/107
Hidden layer sizes 64/64
Activation Function ReLU
Loss Function MAE
Optimizer Adam
Weights initialization Xavier
Learning rate $\alpha$ $8\cdot 10^{3}$
Epochs 5000
Regularization Dropout, 0.001
Mini batch Size 128
Scheduler: $\gamma$, Step 0.975, 30
Appendix C Statistical validation of moment closure and finite sample deviations
As stated in Section 2, our method predicates on approximating the individual damages in each flight –the latent random variables $\mathscr{D}_{\text{GAG}}$ and $\mathscr{D}_{\text{G\&M}}$ – by their sample mean. In this appendix, we give insight into the validity of the implicit assumptions behind this procedure with an analysis based on the GAG damage.
There are mainly two approximations:
  • 1.
    First, we identify each realization of the GAG damage random variable $\mathscr{D}_{\text{GAG}}$ with the observed per-flight average $\overline{\mathscr{D}_{\text{GAG}}}$, i.e, we approximate a random variable with its (estimated) first moment. This is sometimes called a mean-field approximation or a first-moment closure, and it is a valid approximation when higher moments of the distribution (e.g. fluctuations of the random variable around its mean value) are relatively small (in particular, much smaller than the mean). Such comparison can be made by computing the so-called squared coefficient of variation, $\text{CV}^{2}=\frac{\sigma^{2}}{\overline{\mathscr{D}_{\text{GAG}}}^{2}}$, where $\sigma^{2}$ and $\overline{\mathscr{D}_{\text{GAG}}}$ are the sample variance and sample mean of the GAG damage per-flight, respectively. First moment closures usually require $\text{CV}^{2}\ll 1$. In Table 15 we provide an illustrative table with results for three missions for a given PSE and $k_{t}$. As observed, identifying each realization of the per-flight random variable $\mathscr{D}_{\text{GAG}}$ with its sample average is not fully justified yet, $\text{CV}^{2}\leq 1$, not $\text{CV}^{2}\ll 1$, suggesting non-negligible fluctuations. However, as specified in section 2 the damage of one particular flight is not the variable of interest. On the contrary, the important variable is the cumulated damage after $n$ flights, $\mathbb{D}_{\text{GAG}}=\sum_{i=1}^{n}\mathscr{D}_{\text{GAG}}$. Accordingly, the first moment closure needs to hold with respect to this cumulated random variable $\mathbb{D}_{\text{GAG}}$. Now, for large $n$ (when expected values coincide with sample averages), the expected value of the cumulated GAG damage $\mathbb{E}(\mathbb{D}_{\text{GAG}})\equiv\overline{\mathbb{D}_{\text{GAG}}}$, which for ease of notation we label it $\mathbb{D}_{\text{GAG}}$, reads $\mathbb{D}_{\text{GAG}}=\overline{\mathscr{D}_{\text{GAG}}}\cdot n$, since
    $\mathbb{E}(\mathbb{D}_{\text{GAG}})=\mathbb{E}(\sum_{i=1}^{n}\mathscr{D}_{\text{GAG}})=\sum_{i=1}^{n}\mathbb{E}(\mathscr{D}_{\text{GAG}})=n\overline{\mathscr{D}_{\text{GAG}}}.$ (7)
    At the same time, as we are assuming flight independence we have
    $\text{Var}(\mathbb{D}_{\text{GAG}})=\text{Var}(\sum_{i=1}^{n}\mathscr{D}_{\text{GAG}})=\sum_{i=1}^{n}\text{Var}(\mathscr{D}_{\text{GAG}})=n\cdot\sigma^{2}.$ (8)
    Accordingly, assessment of the first moment closure approximation via the squared coefficient of variation reads now
    $\text{CV}^{2}_{\mathbb{D}_{\text{GAG}}}=\frac{\text{Var}(\mathbb{D}_{\text{GAG}})}{\mathbb{E}(\mathbb{D}_{\text{GAG}})^{2}}=\frac{n\cdot\sigma^{2}}{n^{2}\cdot\overline{\mathscr{D}_{\text{GAG}}}^{2}}=\frac{\text{CV}^{2}}{n}$ (9)
    In short, the first moment closure is a substantially improved approximation for large $n$ (see Table 15 for three different missions), and overall the procedure is thus sound and justified.
    $n$ $\text{CV}^{2}$ $\text{CV}^{2}_{\mathbb{D}_{\text{GAG}}}$
    $\approx 800$ 1 0.001
    $\approx 1300$ 0.3 0.0002
    $\approx 13000$ 0.09 0.000007
    Table 15: Squared coefficients of variation for $\mathscr{D}_{\text{GAG}}$ and $\mathbb{D}_{\text{GAG}}$, for three different PSEs illustrating very different number of flights $n$.
  • 2.
    Second, approximating the (true, hidden) expected value of a random variable by its sample mean is a classical estimation method, which holds by virtue of the law of large numbers which assures almost-surely convergence of the sample mean to the true mean for large sample sizes. For illustration, we assess how the (true, hidden) average per-flight GAG damage deviates from the (estimated) sample average per-flight GAG damage $\overline{\mathscr{D}_{\text{GAG}}}$ using non-parametric bootstrapping and building the 95% confidence interval of the bootstrapped mean GAG Damage, as this gives us an uncertainty interval where the true mean will lies with high probability. In Figure 18 we represent, for a specific PSE, $k_{t}$ and mission ($n\approx 1300$ flights), the distribution of bootstrapped means alongside its 2.5 and 97.5 percentiles as well as the bootstrap average. We also locate the original sample mean for reference. Observe that the sample mean of the GAG damage, $\overline{\mathscr{D}_{\text{GAG}}}$ and the average of the bootstrapped means $\mu$ virtually coincide, $\overline{\mathscr{D}_{\text{GAG}}}\approx\mu=6.19\times 10^{-5}$, and the bootstrap $CI95=[6.10\times 10^{-5},6.29\times 10^{-5}]$ is a very narrow one. This validates our procedure of identifying the true mean with the sample mean. Results hold similarly for other cases.
Figure 18: Distribution of bootstrapped means associated to the GAG damage $\mathscr{D}_{\text{GAG}}$, for a specific PSE and mission. The 2.5 and 97.5 percentiles –conforming the bootstrap CI95 of average per-flight GAG damage are highlighted, alongside the average of the bootstrapped means $\mu$ and the sample mean $\overline{\mathscr{D}_{\text{GAG}}}$.
This procedure critically depends on the number of flights $n$ being moderately large, something that in this work systematically holds as we have $n\in[800,13000]$. Now, it is always desirable to establish what would be the minimum number of flights $n_{\text{min}}$ such that our method is acceptable. This is akin to ask for what sample size (i) the first moment closure is no longer a valid one, or (ii) when estimating the true mean with the sample mean is not acceptable. The actual question is difficult to answer in general, as it requires to define what we mean by ‘acceptable’. Here we provide two different roadmaps to estimate the minimum $n_{min}$ to meet a certain statistical guarantee, namely, that with a certain significance level $\alpha$ and confidence level $p=1-\alpha$, the mismatch between the sample mean of the per-flight GAG damage $\overline{\mathscr{D}_{\text{GAG}}}$ and the true mean is smaller than a certain tolerance $\epsilon$ (measured in relative error units).
In general terms, invoking the central limit theorem (CLT, which holds as high order moments of $\mathscr{D}_{\text{GAG}}$ do not diverge, see above), such difference drops to zero as fast as $1/\sqrt{(}N)$. In the event that $\mathscr{D}_{\text{GAG}}$ is Gaussian, simple probability arguments provide a closed expression for $n_{min}$:
$n_{min}=\left(\frac{Z(\alpha/2)\sigma}{\epsilon\overline{\mathscr{D}_{\text{GAG}}}}\right)^{2},$ (10)
where $\alpha$ is our desired significance level, $Z(\alpha/2)$ can be obtained from the values from a standard normal distribution, and sample standard deviation $\sigma$ and the sample mean $\overline{\mathscr{D}_{\text{GAG}}}$ can be computed from the population dataset. When the marginal distribution deviates from a pure Gaussian, the expression above is only an approximation, which nonetheless becomes a tight one as soon as $n$ is moderately large (our case) and if the underlying distribution does not have strong fluctuations around the mean (i.e. long tails), which, again, is the case here.
For illustration, for the concrete mission and PSE discussed above with $n\approx 1300$, and fixing $\epsilon=0.02$ and $\alpha=0.05$ (i.e. allowing for a maximum 2% relative error with a 95% confidence), we find $n_{min}$ = 883, which is substantially smaller than the actual number of flights. Similarly, we can estimate the maximum error associated with the exact $n$ used. In this case, for a 95% confidence level, the maximum error is $1.64\%$.
Finally, we can also define an heuristic which makes use of nonparametric bootstrapping (and therefore does not assume any shape for the GAG Damage distribution). The minimal acceptable size of the dataset can be found by (i) establishing an acceptable interval for the sample mean in the original dataset (e.g. the CI95 of the bootstrapped mean), then (ii) sequentially downsampling the original dataset, estimating the new sample mean in each case, and (iii) detecting when the sample mean is no longer within the established interval.
In practice, the procedure is implemented as follows: first, we define a new data size $n_{new}<n$. Then, we perform independent resamplings (here we performed 2000, with replacement) of the original dataset of size $n_{new}$, and compute the percentage of cases where the new sample mean of the resampled datasets falls within the originally established interval.
These ‘acceptable intervals’ were built with two different approaches, each providing different and complementary information: i) the interval is defined by fixing a priori a relative error around the average of the bootstrapped means $\mu$ (see e.g. Figure 18) so as to build the interval $\mu-\epsilon\mu,\mu+\epsilon\mu]$, and ii) the interval is defined as the 95% confidence interval of the bootstrapped mean.
For illustration, using the same mission and PSE with $n\approx 1300$, in Figure 19 we show the percentage of subsets falling within each interval respectively, as a function of $n_{new}.$ Observe that for a sample size of 883 (flagged as $n_{\text{min}}$ by the analytical approach above), in about $95\%$ of the resamples the sample average error is equal or lower than the established error, (2%), whereas $87.5\%$ of the samples fall in the CI95 interval.
Figure 19: (Left panel) Percentage of resamples with $n_{\text{new}}<n$ data where the sample mean is within the acceptable interval defined initially as $\mu-\epsilon\mu,\mu+\epsilon\mu]$, as a function of the datasize sample $n_{\text{new}}$. (Right panel) Similar to the left panel, but here the acceptable interval is defined as the bootstrap mean CI95.